OpenStudy (anonymous):

insert A.M 's between 7 and 71 such that the fifth A.M is 27 Find the number of A.M's

OpenStudy (experimentx):

how do you insert AM's in 1,_,_,a ?

OpenStudy (anonymous):

(a-1)/ (m+1) = d

OpenStudy (anonymous):

1 +d , 1 + 2d.........

OpenStudy (anonymous):

@experimentX there was a mistake in my question....plzz...read it nw

OpenStudy (experimentx):

let there be n AM's between 7 and 71 71 = 7 + nd 27 = 7 + 4d shouldn't be this way ... seems i forgot too. time to google for AM's http://en.wikipedia.org/wiki/Arithmetic_mean

OpenStudy (anonymous):

But that Didnt work

OpenStudy (anonymous):

27 = 7 + 4d d = 5 71 = 7 + nd 64 = 5n n = 64/5,..............

OpenStudy (experimentx):

of course this doesn't work because i seem to have forgotten the formula for AM's

OpenStudy (anonymous):

A.M a , b , c b = (a + c)/2

OpenStudy (experimentx):

i know that ... a, _, _, b

OpenStudy (experimentx):

a, x, y, b x = ((a+b)/2 + a)/2 y = (x+b)/2

OpenStudy (experimentx):

was A.S. after all http://www.wolframalpha.com/input/?i=solve+71+%3D+7+%2B+n+d%2C+27+%3D+7+%2B+5d

OpenStudy (anonymous):

Foolish.....)...i made a fatal....Mistake...in my calculation...hm.....thxxxxx......))